# Learning a Gaussian Mixture for Sparsity Regularization in Inverse Problems

### v1

@@ Abstract.
To evaluate the effectiveness of our approach, we conduct a numerical comparison with commonly employed sparsity-promoting regularization techniques, namely LASSO, group LASSO, iterative hard thresholding and sparse coding/dictionary learning. Notably, our reconstructions consistently exhibit lower mean square error values across all $1$D datasets utilized for the comparisons, even in cases where the datasets significantly deviate from a Gaussian mixture model.
@@ 1. Introduction
The key ingredient needed in sparse optimization is the knowledge of a suitable dictionary that can well represent the unknown quantities with as few elements as possible. If these quantities cannot be described analytically and are measured from experiments, given a sufficient number of samples, machine learning techniques can be used to infer the dictionary [17]. Dictionary learning in the context of sparse optimization, i.e. sparse coding, can be seen as the problem of finding the best change of basis that makes the data as sparse as possible. Thus, it is a special case of the problem of learning a regularization functional [20, 26, 1] and the more general framework of learning operators between function spaces (see [15] and references therein).
@@
Our main contribution is the development of supervised and unsupervised learning schemes to approximate the Bayes estimator. This is done thanks to an explicit expression for the MMSE, which is usually not available for general probability distributions underlying the data and the noise. The proposed training schemes are based on the observation that the explicit expression of the MMSE can be seen as a two-layer neural network, which can be easily trained using stochastic optimization and back-propagation. In particular, we find a strong similarity of our network with a single self-attention layer, a building block of the well-known transformer architecture [29]: our results provide a novel statistical interpretation of the attention mechanism.
After the training, our network/algorithm can be used as an alternative to classical sparse optimization approaches. We perform extensive numerical comparisons of our approach with well-known sparsity-promoting algorithms: LASSO, group LASSO, iterative hard thresholding (IHT), and sparse coding/dictionary learning. The tests are done by performing denoising and deblurring on several 1D datasets. In all experiments, our method outperforms the others, in the MSE. Our method can be seen as a new paradigm for both sparse optimization and dictionary learning for sparse data.
@@
The paper is organized as follows. In Section 2, we provide an overview of the concept of Bayes estimator and recall an explicit formula in the case of linear inverse problems with Gaussian mixture prior. Afterwards, Section 3 is dedicated to the interpretation of this estimator as a neural network, highlighting its potential utility for sparse recovery. We introduce a supervised and an unsupervised algorithm for learning the neural network representation of the Bayes estimator in Section 4. In Section 5, we describe some baseline sparsity-promoting algorithms, against which our methods are evaluated. This is done in Section 6, where we conduct numerical comparisons on denoising and deblurring problems across various 1D datasets. Conclusions are drawn in Section 7. Further contents are reported in the appendices, including the proof of the main theoretical result (Appendix A), implementation details about the baseline algorithms (Appendix B), and extended numerical comparisons (Appendix C).
@@ 2. Statistical learning and Bayes estimator for inverse problems
In this context, it is easy to construct a hypothesis class $\mathcal{H}$ that contains the Bayes estimators $R^{\star}$ of every possible prior in the parametric class, namely,
@@
where $\rho(\theta)$ is the joint probability distribution for $(X,Y)$ when $X\sim\rho_{X}(\theta)$, $Y=AX+E$ and $E\sim\rho_{E}$. Recall that, by Assumption 2.1, the distribution of the noise $E$ is known, and so it is considered fixed.
@@
Despite this approach being rather general, it is useful only if it is possible to provide a closed-form expression of $R_{\theta}$, which might entail a learning approach for its approximation. This can be easily done when $X$ belongs to fairly simple classes, such as Gaussian random variables or, as we show in this section, mixtures of Gaussians.
@@ 2.1. Bayes estimator for linear inverse problems with a Gaussian Mixture prior
Let $X$ be a Gaussian mixture in $\mathbb{R}^{n}$, as in Definition  2.3, and $E\sim\mathcal{N}(\mathbf{0},\Sigma_{E})$ be independent of $X_{i}$ for every $i$, and of $I$. Let $A\in\mathbb{R}^{m\times n}$ be such that $A\Sigma_{i}A^{T}+\Sigma_{E}$ is invertible for every $i=1,\dots,L$. Set $Y=AX+E$, as in (3). The corresponding Bayes estimator is
@@
A proof of Theorem 2.4 can be found in [16]. For completeness, we provide a similar proof in Appendix A.
@@ 3.1. Bayes estimator as a neural network
collecting all the weights, means, and covariances of the mixture. We can now denote the function defined in (5) as $R_{\theta}$, parametrized according to (7). Note that the forward map $A$ and the covariance of the noise $\Sigma_{E}$ are considered known and fixed, and we assume that $\Sigma_{E}$ is invertible, so that $A\Sigma_{i}A^{T}+\Sigma_{E}$ is invertible for every $i$. The resulting hypothesis class is
@@
We now focus on describing formula (5) in terms of a neural network’s architecture (see Figure 1):
@@
The transformer architecture, based on the attention mechanism [29], has revolutionized the field of machine learning, especially natural language processing (NLP), offering a powerful solution for capturing and utilizing relevant information within complex data. Inspired by human attention, this mechanism enables models to selectively focus on the most important elements of the input, significantly enhancing their performance across a wide range of tasks, from NLP to computer vision.
Traditionally, machine learning models treat all input elements equally, neglecting the varying importance of different parts of the data. Attention addresses this limitation by introducing a mechanism that assigns weights or attention scores to different elements. By doing so, models can dynamically allocate their resources to the most relevant components, so their interpretability and performance increase.
@@
(12) $\text{Att}(\eta)=\text{Att}(Q,K,V):=\mathrm{softmax}\bigg{(}\frac{QK^{T}}{\sqrt{L}}\bigg{)}V\in\mathbb{R}^{L\times m},$
@@
Considering as input $\eta\in\mathbb{R}^{L\times m}$ with rows $\eta_{i}:=y-A\mu_{i}\in\mathbb{R}^{m}$, $i=1,...,L$, we are able to define three affine maps (not linear as in the classical attention mechanism) modeling the queries, the keys and the values. More precisely, the rows of $Q$ and $K$ are, respectively,
$Q_{i}=\frac{l_{i}^{\frac{1}{2}}}{\sqrt{m}}\mathbf{1}_{m}+\frac{1}{\sqrt{2}}M_{i}\eta_{i}\in\mathbb{R}^{m}$
and
@@
where $l_{i}=\log{\bigg{(}\frac{w_{i}}{\sqrt{(2\pi)^{n}|A\Sigma_{i}A^{T}+\Sigma_{E}|}}\bigg{)}}$, $M_{i}=(A\Sigma_{i}A^{T}+\Sigma_{E})^{-\frac{1}{2}}$ and $\mathbf{1}_{m}=(1,\dots,1)\in\mathbb{R}^{m}$. Then $z_{i}\in\mathbb{R}$, $i=1,...,L$, defined in (10) can be rewritten as
$z=(Q\odot K)\mathbf{1}_{m}^{T}\in\mathbb{R}^{L\times 1},$
where $\odot$ represents the Hadamard product (or element-wise product).
Next, by defining the rows of $V\in\mathbb{R}^{L\times n}$ where $n$ is the dimension of the linear inverse problem unknowns (1) using the affine transformation of $\eta$ given in (11), namely
@@
we obtain
(13) $R^{\star}(y)=\overline{\text{Att}}(\eta)=\overline{\text{Att}}(Q,K,V):=\mathrm{softmax}\bigg{(}(Q\odot K)\mathbf{1}_{m}^{T}\bigg{)}^{T}V\in\mathbb{R}^{n}.$
@@
While Gaussian distributions work well to represent smoothness priors, they are not well-adapted to model sparsity. Here, we propose to use (degenerate) Gaussian mixture models to represent (group) sparsity prior. An alternative approach using hierarchical Bayesian models is considered in [5, 6]. We also note that our unsupervised approach (introduced below) shares a similar clustering step with the dictionary learning strategy presented in [4] within a hierarchical model framework.
@@ 4. Proposed algorithms: Supervised and Unsupervised approaches
In this section, we propose two different algorithms to learn the neural network representation (9) of the Bayes estimator (5) to solve the statistical linear inverse problem of retrieving $X$ from $Y$ in (3). We assume that the random variables $X$ and $E$ satisfy Assumption 2.1, and write $X$ as a mixture of $L$ random variables
@@ 4.1. Supervised approach
where $\lambda$ is a regularization parameter, $\mathcal{J}(\theta)=\sum_{i=1}^{L}\|\Sigma_{i}\|_{*}$, and $*\|\cdot\|_{*}$ is the nuclear norm. Such a regularization term can be equivalently defined as the $\ell^{1}$ norm of the singular values of $\Sigma_{i}$, hence it promotes low-rank covariances without requiring the knowledge of their eigenvectors.
The evaluation of the nuclear norm is computationally expensive and should be done $L$ times for each step of the minimization process. Another option is taking $\mathcal{J}(\theta)=\sum_{i=1}^{L}\|\Sigma_{i}\|^{2}_{\text{F}}$, where $\|\cdot\|_{\text{F}}$ represents the Frobenius norm. This choice is computationally convenient, but does not promote any kind of sparsity on the covariances. However, in the numerical simulations, we do not observe significant differences when using the nuclear norm, the Frobenius norm or no regularization term. Therefore, in the numerical results of Section 6 we did not make use of any regularization term.
@@ 5.2. Group LASSO
where $\beta_{i}$ is the coefficient vector corresponding to the $i$-th group, $\beta\in\mathbb{R}^{nL}$ is the concatenation of the vectors $\beta_{1},\ldots,\beta_{L}$, and $\lambda>0$ is a regularization parameter. The second term of (18) is a regularization functional encouraging entire groups of features to be either included or excluded from the model. The weighted norm $\|\beta_{i}\|_{2}$ with $\|\beta_{i}\|_{K_{i}}:=(\beta_{i}^{T}K_{i}\beta_{i})^{\frac{1}{2}}$ is employed to promote prior information regarding the sparse representation of each group. The minimization of (18) can be performed via the iterative algorithm proposed in [30, Proposition $1$], as discussed in Appendix B. As for LASSO, also for Group LASSO one has to choose the regularization parameter $\lambda$, and to know a priori the group bases $\mathcal{B}_{i}$ for $i=1,...,L$.
@@ 5.4. Dictionary Learning
where $\lambda>0$ is a regularization parameter and $\beta_{j}\in\mathbb{R}^{d}$ is the sparse representation of $x_{j}$ with respect to the dictionary $D$. If the dictionary is fixed, the functional in (20) resembles the LASSO functional (17), except for the fact that the dictionary may not be a basis and that we are summing over the elements of the training set. However, in this case, the goal is mainly to learn $D$. In order to solve the minimization problem in (20), we rely on the online method used in [22], which iteratively minimizes over one variable keeping the others fixed. We update the sparse representation $\beta$ by using the least angle regression (LARS) algorithm and the dictionary $D$ by using block coordinate descent.
@@ 6. Numerical results
In this section, we compare the algorithms proposed in Section 4 with the baseline methods discussed in Section 5 for $1$D denoising and deblurring problems with three datasets. Our experiments primarily aim to compare our algorithms with classical sparsity-promoting methods on simplified signal classes, rather than striving for state-of-the-art results. The exploration of more complex inverse problems and the incorporation of real-world data fall outside the scope of this paper and are designated for future research.
@@ 6.2. Methods
In the following experiments, we test our supervised and unsupervised algorithms described in Sections 4.1 and 4.2 and compare them with the baseline algorithms described in Section 5. More precisely, we consider the following methods.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract.
To evaluate the effectiveness of our approach, we conduct a numerical comparison with commonly employed sparsity-promoting regularization techniques, namely LASSO, group LASSO, iterative hard thresholding, and sparse coding/dictionary learning. Notably, our reconstructions consistently exhibit lower mean square error values across all $1$D datasets utilized for the comparisons, even in cases where the datasets significantly deviate from a Gaussian mixture model.
@@ 1. Introduction
The key ingredient needed in sparse optimization is the knowledge of a suitable dictionary that can well represent the unknown quantities with as few elements as possible. If these quantities cannot be described analytically and are measured from experiments, given a sufficient number of samples, machine learning techniques can be used to infer the dictionary [21, 17, 18]. Dictionary learning in the context of sparse optimization, i.e. sparse coding, can be seen as the problem of finding the best change of basis that makes the data as sparse as possible. Thus, it is a special case of the problem of learning a regularization functional [24, 30, 1] and the more general framework of learning operators between function spaces (see [10, 19] and references therein).
@@
Our main contribution is the development of supervised and unsupervised learning schemes to approximate the Bayes estimator when a (degenerate) Gaussian Mixture model is assumed as a prior on $X$. This is done thanks to an explicit expression for the MMSE, which is usually not available for general probability distributions underlying the data and the noise. The proposed training schemes are based on the observation that the explicit expression of the MMSE can be seen as a two-layer neural network, which can be easily trained using stochastic optimization and back-propagation. In particular, we find a strong similarity of our network with a single self-attention layer, a building block of the well-known transformer architecture [33]: our results provide a novel statistical interpretation of the attention mechanism.
As carefully presented in Section 4, the supervised technique performs an empirical risk minimization within the hypothesis class $\mathcal{H}$ of Bayes estimators of Gaussian Mixture models, the means and the covariance matrices of the mixture being learnable parameters. The unsupervised one, instead, approximates the Bayes estimator relying on suitable clustering techniques and empirical estimates for the means and covariances.
After the training, our network/algorithm can be used as an alternative to classical sparse optimization approaches. We perform extensive numerical comparisons of our approach with well-known sparsity-promoting algorithms: LASSO, group LASSO, iterative hard thresholding (IHT), and sparse coding/dictionary learning. The tests are done by performing denoising and deblurring on several datasets of (discretized) 1D signals. In all experiments, our method outperforms the others, in the MSE. Our method can be seen as a new paradigm for both sparse optimization and dictionary learning for sparse data.
@@
The paper is organized as follows. In Section 2, we provide an overview of the concept of Bayes estimator and recall an explicit formula in the case of linear inverse problems with Gaussian mixture prior. Afterwards, Section 3 is dedicated to the interpretation of this estimator as a neural network, highlighting its potential utility for sparse recovery. We introduce a supervised and an unsupervised approach for learning the neural network representation of the Bayes estimator in Section 4. In Section 5, we describe some baseline sparsity-promoting algorithms, against which our methods are evaluated. This is done in Section 6, where we conduct numerical comparisons on denoising and deblurring problems across various datasets of 1D signals. Conclusions are drawn in Section 7. Further contents are reported in the appendices, including the proof of the main theoretical result (Appendix A), implementation details about the baseline algorithms (Appendix B), and extended numerical comparisons (Appendix C).
@@ 2. Statistical learning and Bayes estimator for inverse problems
In this context, it is natural to construct a hypothesis class $\mathcal{H}$ that contains the Bayes estimators $R^{\star}$ of every possible prior in the parametric class, namely,
@@
where $\rho(\theta)$ is the joint probability distribution for $(X,Y)$ when $X\sim\rho_{X}(\theta)$, $Y=AX+E$ and $E\sim\rho_{E}$. Recall that, by Assumption 2.1, the distribution of the noise $E$ is known, and so it is considered fixed. However, the choice of this hypothesis class is non-standard, if compared to deep learning approaches exploiting the approximation power of neural networks. The two main motivations for its adoption are that the Bayes estimator has sharp statistical properties in terms of MSE, and moreover it is well-suited for both supervised and unsupervised learning approaches, as we explain below.
@@
Despite this approach being rather general, it is useful only if it is possible to provide a closed-form expression of $R_{\theta}$, which might entail a learning approach for its approximation. This can be easily done when $X$ belongs to fairly simple classes, such as Gaussian random variables or, as we show in this section, mixtures of Gaussians. Other possible examples entailing closed-form representation of the posterior distribution (thus, of the Bayes estimator) are represented by the families of conjugate priors associated with the specified likelihood, i.e., with the noise model: for example, in the case of a Poisson likelihood, if the prior has a Gamma distribution, then also the posterior does.
@@ 2.1. Bayes estimator for linear inverse problems with a Gaussian Mixture prior
Let $X$ be a Gaussian mixture in $\mathbb{R}^{n}$, as in Definition  2.3, and let $E\sim\mathcal{N}(\mathbf{0},\Sigma_{E})$ be such that Assumption 2.1 is verified, i.e., $\Sigma_{E}$ is invertible and $E$ is independent of $X_{i}$ for every $i$, and of $I$. Set $Y=AX+E$, as in (3). The corresponding Bayes estimator is
@@
Note that the forward map $A$ and the covariance of the noise $\Sigma_{E}$ are considered known and fixed, and we assume that $\Sigma_{E}$ is invertible, so that $A\Sigma_{i}A^{T}+\Sigma_{E}$ is invertible for every $i$. A proof of Theorem 2.4 can be found in [20]. For completeness, we provide a similar proof in Appendix A.
@@ 3.1. Bayes estimator as a neural network
collecting all the weights, means, and covariances of the mixture. We can now denote the function defined in (5) as $R_{\theta}$, parametrized according to (7). The resulting hypothesis class is
@@
We now focus on describing formula (5) in terms of a neural network’s architecture (see Figure 1). We have the following equivalent formula for the Bayes estimator.
Proposition 3.1.
We have that
@@
The transformer architecture, which uses the attention mechanism [33], has greatly improved machine learning, especially in natural language processing (NLP). This mechanism helps models focus on the most important parts of the input, boosting performance in various tasks, including NLP and computer vision. Instead of treating all input elements the same, the attention mechanism gives different weights to different parts, allowing models to allocate resources better and enhance both interpretability and performance.
@@
(12) $\operatorname{Att}(\eta)=\operatorname{Att}(Q,K,V):=\mathrm{softmax}\bigg{(}\frac{QK^{T}}{\sqrt{L}}\bigg{)}V\in\mathbb{R}^{L\times m},$
@@
In our case, we have the following result, as an immediate consequence of Proposition 3.1.
Proposition 3.2.
The Bayes estimator for a Gaussian mixture can be written as follows:
(13) $R^{\star}(y)=\overline{\operatorname{Att}}(\eta)=\overline{\operatorname{Att}}(Q,K,V):=\mathrm{softmax}\bigg{(}(Q\odot K)\mathbf{1}_{m}^{T}\bigg{)}^{T}V\in\mathbb{R}^{n},$
where $\odot$ represents the Hadamard product (or element-wise product), the matrices $Q,K,V\in\mathbb{R}^{L\times m}$ have rows defined, for $i=1,...,L$, by
  • $Q_{i}=\frac{l_{i}^{\frac{1}{2}}}{\sqrt{m}}\mathbf{1}_{m}+\frac{1}{\sqrt{2}}M_{i}\eta_{i}\in\mathbb{R}^{m}$,
@@
where
  • $\eta_{i}=y-A\mu_{i}\in\mathbb{R}^{m}$,
  • $l_{i}=\log{\bigg{(}\frac{w_{i}}{\sqrt{(2\pi)^{n}|A\Sigma_{i}A^{T}+\Sigma_{E}|}}\bigg{)}}$,
  • $M_{i}=(A\Sigma_{i}A^{T}+\Sigma_{E})^{-\frac{1}{2}}$,
  • $\mathbf{1}_{m}=(1,\dots,1)\in\mathbb{R}^{m}$.
@@
While Gaussian distributions work well to represent smoothness priors, they are not well-adapted to model sparsity. Here, we propose to use (degenerate) Gaussian mixture models to represent (group) sparsity prior. In the context of imaging inverse problems, the use of Gaussian Mixture Models to describe structured sparsity has been leveraged in [34], although focusing on maximum-a-posteriori rather than Bayes estimation. An alternative approach using hierarchical Bayesian models is considered in [5, 6]. We also note that our unsupervised approach (introduced below) shares a similar clustering step with the dictionary learning strategy presented in [4] within a hierarchical model framework.
@@ 4. Proposed algorithms: Supervised and Unsupervised approaches
In this section, we propose two different techniques to learn the neural network representation (9) of the Bayes estimator (5) to solve the statistical linear inverse problem of retrieving $X$ from $Y$ in (3). We assume that the random variables $X$ and $E$ satisfy Assumption 2.1, and write $X$ as a mixture of $L$ random variables
@@ 4.1. Supervised approach
where $\lambda$ is a regularization parameter, $\mathcal{J}(\theta)=\sum_{i=1}^{L}\|\Sigma_{i}\|_{*}$, and $\|\cdot\|_{*}$ is the nuclear norm. Such a regularization term can be equivalently defined as the $\ell^{1}$ norm of the singular values of $\Sigma_{i}$, hence it promotes low-rank covariances without requiring the knowledge of their eigenvectors.
The evaluation of the nuclear norm is computationally expensive and should be done $L$ times for each step of the minimization process. Another option is taking $\mathcal{J}(\theta)=\sum_{i=1}^{L}\|\Sigma_{i}\|^{2}_{\text{F}}$, where $\|\cdot\|_{\text{F}}$ represents the Frobenius norm. This choice is computationally convenient, but does not promote any kind of sparsity on the covariances. However, in the numerical simulations, we do not observe significant differences when using the nuclear norm, the Frobenius norm or no regularization term (see in particular Table 3). Therefore, in the numerical results of Section 6 we did not make use of any regularization term.
@@ 5.2. Group LASSO
where $\beta_{i}$ is the coefficient vector corresponding to the $i$-th group, $\beta\in\mathbb{R}^{nL}$ is the concatenation of the vectors $\beta_{1},\ldots,\beta_{L}$, and $\lambda>0$ is a regularization parameter. The second term of (18) is a regularization functional encouraging entire groups of features to be either included or excluded from the model. The weighted norm $\|\beta_{i}\|_{2}$ with $\|\beta_{i}\|_{K_{i}}:=(\beta_{i}^{T}K_{i}\beta_{i})^{\frac{1}{2}}$ is employed to promote prior information regarding the sparse representation of each group. The minimization of (18) can be performed via the iterative algorithm proposed in [35, Proposition $1$], as discussed in Appendix B. As for LASSO, also for Group LASSO one has to choose the regularization parameter $\lambda$, and to know a priori the group bases $\mathcal{B}_{i}$ for $i=1,...,L$.
@@ 5.4. Dictionary Learning
where $\lambda>0$ is a regularization parameter and $\beta_{j}\in\mathbb{R}^{d}$ is the sparse representation of $x_{j}$ with respect to the dictionary $D$. If the dictionary is fixed, the functional in (20) resembles the LASSO functional (17), except for the fact that the dictionary may not be a basis and that we are summing over the elements of the training set. However, in this case, the goal is mainly to learn $D$. In order to solve the minimization problem in (20), we rely on the online method used in [26], which iteratively minimizes over one variable keeping the other ones fixed. We update the sparse representation $\beta$ using the least angle regression algorithm (LARS) and the dictionary $D$ by block coordinate descent. As reported in [26], the computational cost of each iteration of the online method is dominated by the update of $\beta$, and the complexity of LARS applied on $D\in\mathbb{R}^{n\times d}$ is of order $d^{3}+d^{2}n$ (see [12]).
@@ 6. Numerical results
In this section, we compare the methods proposed in Section 4 with the baseline algorithms discussed in Section 5 for $1$D denoising and deblurring problems with three datasets. Our experiments primarily aim to compare our methods with classical sparsity-promoting techniques on simplified signal classes, rather than striving for state-of-the-art results. The exploration of more complex inverse problems and the incorporation of real-world data fall outside the scope of this paper and are designated for future research.
@@ 6.2. Methods
In the following experiments, we test our supervised and unsupervised approaches described in Sections 4.1 and 4.2 and compare them with the baseline algorithms described in Section 5. More precisely, we consider the following methods.
@@
    Supervised(§$\S$4.1). We consider both the case in which the weights of the network are randomly initialized, denoted as (A), and the one in which they are initialized by the values obtained as an outcome of the unsupervised procedure, denoted as (A+B).
@@
    Dictionary learning (§$\S$5.4). We learn a sparsifying dictionary $D$ for the training set, using the Lasso LARS algorithm to solve (20); then, we employ the learned $D$ to denoise the test signal by minimizing (21). We choose the number of elements of the dictionary as $d=\frac{n}{2}$, where $n$ is the signal length, in order to balance expressivity and numerical efficiency.
@@
    Group dictionary learning (§$\S$5.4). After dividing the training set into $L$ clusters and learning a sparsifying dictionary $D_{i}$ for each group, we minimize (22). In this case, we choose $d=\frac{n}{2L}$.
@@
    IHT with SVD basis (§$\S$5.3). We infer the basis with respect to which the unknown is sparse by computing the SVD of the empirical covariance matrix of the training set and by considering the orthonormal basis composed by the eigenvectors. Then, we choose $\mathcal{S}$ as the union of all the $s$-dimensional coordinate subspaces in $\mathbb{R}^{n}$ w.r.t. the inferred sparsity basis. Here we choose the degree of sparsity $s$ by minimizing the relative MSE (computed w.r.t. the norm of the original signals) over the training set.
@@
    IHT with SVD bases of groups (§$\S$5.3). We divide the training set into $L$ clusters as explained in Section 4.2. Then, for each group provided by the clustering, we infer the basis with respect to which that group is sparse by computing the SVD of its empirical covariance matrix and by considering the eigenvectors corresponding to the $s$ largest eigenvalues, where the degree of sparsity $s$ is chosen as for (E). Finally, we choose $\mathcal{S}$ as the union of the $s$-dimensional subspaces spanned by the bases inferred.
@@
    IHT with known basis (§$\S$5.3). We suppose to know the basis with respect to which the unknown is sparse. In particular, for Dataset 1 we consider the canonical basis, while Datasets 2 and 3 require a more complicated treatment. Indeed, as discussed in Section 6.1, their representation in a suitable wavelet basis is sparse at fine scales. For this reason, we preprocess the datasets by applying the wavelet transform, keep the low-scale coefficients fixed, and apply IHT only to the fine scales. We choose the wavelet basis generated by the Daubechies wavelet with $6$ vanishing moments.
@@
    LASSO with SVD basis (§$\S$5.1). We infer the basis with respect to which the unknown is sparse as in (E). Then, we minimize (17).
@@
    Group LASSO with SVD bases (§$\S$5.2). We divide the training set into $L$ clusters as explained in Section 4.2. Then, for each group provided by the clustering, we infer the basis with respect to which that group is sparse as explained in (F). Here, however, we consider the complete bases provided by the SVDs of the empirical covariance matrices of the groups. We choose the power $-\frac{1}{2}$ of the empirical covariance matrices of each group as penalty matrices $K_{i}$.
@@
    LASSO with known basis (§$\S$5.1). We incorporate in LASSO the knowledge of the basis with respect to which the unknown is sparse, as already explained for (G). This implies that, for Dataset 1, we solve LASSO with the canonical basis, whereas for Datasets 2 and 3 we first compute the wavelet transform and then solve LASSO only on the high-scale coefficients.